OpenStudy (eric_d):

cos^4 theta - sin^4 theta divide by cos theta-sin theta

OpenStudy (eric_d):

@matricked

OpenStudy (anonymous):

a^2 - b^2 =(a+b)(a-b)

OpenStudy (anonymous):

cos^4 theta - sin^4 theta =( cos^2 theta - sin^2 theta )*(cos^2 theta + sin^2 theta ) =( cos^2 theta - sin^2 theta ) [ cos^2 theta + sin^2 theta =1 ] =( cos theta - sin theta ) ( cos theta + sin theta ) i hope the rest u can do

OpenStudy (eric_d):

I can do the rest Thanks @matricked